Using component method of getting the resultant of vectors, find the resultant of the following vectors: 7N along the positive x axis, 5N making an angle of 35o with the x axis, and 3N, 36o South of West.


1
Expert's answer
2021-10-20T10:40:00-0400

Determine the corresponding components of the resultant vector:


"A_x=7\text{ N},\\
A_y=0,\\
B_x=5\cos35°=4.1\text{ N},\\
B_y=5\sin35°=2.9\text{ N},\\
C_x=-3\cos36°=-2.4\text{ N},\\
C_y=-3\sin36°=-1.8\text{ N}.\\\space\\
R_x=A_x+B_x+C_x=9.5\text{ N},\\
R_y=A_y+B_y+C_y=1.1\text{ N},\\
R=\sqrt{R_x^2+R_y^2}=9.6\text{ N},\\\space\\
\theta=\arctan\frac{R_y}{R_x}=6.6°\text{ above x-axis}."